A 79.6 kg man jumps from a window 2.45m above a sidewalk. What is the speed ( in m/s) just before his feet strike the pavement? (m/s)
Answer #1

by newton equation

v^2=u^2+2aS

v^2=0+2*9.8*2.45

v=6.93 m/s